x=3 x=0 directions highschool algebra 1 worksheet

g(x)=3x another problem h(x)=x+7

I donot understand how to do the problem

Answer by checkley71(8403)

G(X)=3X SUBSTITUTE 3 FOR X THUS
G(3)=3*3
G(3)=9
-------------------------
G(0)=3*0
G(0)=0
------------------------
F(X)=X+7
F(3)=3+7
F(3)=10
--------------------------
F(0)=0+7
F(0)=7